OpenStudy (anonymous):

simplify(sin-sin^3)csc

OpenStudy (solomonzelman):

\[(\sin~x-\sin^3x)~\csc~x\]\[(\sin~x-\sin^3x)~\frac{1}{\sin~x}\]\[(1-\sin^2x)\]\[\cos^2x\]